OpenStudy (anonymous):

If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9

OpenStudy (anonymous):

3

mathslover (mathslover):

just remember when you have been given f(x) = x+9 or any thing and also f(x) = 0 then plugin 0 at x f(0) = 0 + 9 = 9 can you do this problem now?